Calculus II, Spring 2013 (http://www.math.nthu.edu.tw/˜ wangwc/)

Homework Assignment for Chap 14

1. Section 14.1: Problems 11, 19, 23, 27, 31, 33, 37, 39, 45.

2. Section 14.3: Problems 9, 13, 15, 23.

3. Section 14.4: Problems 3, 11, 13, 29, 27, 35.

4. Section 14.6: Problems 13, 15, 17, 19, 25, 33, 41, 43.

5. Section 14.7: Problems 3, 5, 8(can you figure out how one chooses the change of variables u, v in this problem?) , 11, 13.

6. Chap 14: 28, 33, 37, 39, 41 (need not evaluate).